# GARLIC: GAussian Representation LearnIng for spaCe partitioning

### v1

@@ 3 Experiments
Evaluation. We evaluate our method on retrieval and classification tasks across the aforementioned datasets. For retrieval, we report Recall10@10 and Recall@1 versus the number of retrieved candidates (probes), reflecting the trade-off between accuracy and efficiency. Recall10@10 measures how many of the true top-10 neighbors are in the retrieved set, while Recall@1 checks if the top-ranked result is the true nearest neighbor. The candidate count controls query time and serves as the main efficiency metric. For MNIST and Fashion-MNIST, we assess classification via majority voting over retrieved candidates: $\hat{y}(q)=\arg\max_{c}\sum_{x\in C_{q}}\mathbb{I}[\text{label}(x)=c]$. This evaluates how well the partitioning preserves semantic similarity.
Baselines and Comparisons. We evaluate our proposed method against several state-of-the-art approximate nearest neighbor search algorithms such as K-Means [6], Neural-LSH (N-LSH) [14], Hyperplane LSH (HP-LSH) [25, 5], Cross-polytope LSH (CP-LSH) [5], ITQ [19], PCA Tree [31, 1, 41, 39], and Faiss-IVF and Faiss-IVFPQFS [15]. For classification tasks, we additionally compare against standard $k$-Nearest Neighbors ($k$-NN) [12, 42], which performs exhaustive search for exact neighbors, and $k$-K-Nearest Neighbors ($kK$-NN) [33], which adjusts the neighborhood size based on local curvature estimates. Each method was evaluated under identical conditions, with hyperparameters optimized according to their original publications to ensure fair and meaningful comparisons across both nearest neighbor retrieval and classification performance.
Table 1: Classification accuracy via majority voting. Ours-1: $\arg\min_{g}\delta_{M}(q,g)$; Ours-2: $\delta_{M}(q,g)\leq\tau$; Ours-3: top-$k$ Gaussians. We highlight the best, second-best and third-best results.
Dataset
Method MNIST $\uparrow$ Fashion-MNIST $\uparrow$
k-NN 0.7576 0.7133
kK-NN 0.7609 0.7796
K-Means 0.9118 0.7589
PCA Tree 0.7750 0.7427
ITQ 0.7156 0.6459
HP-LSH 0.5483 0.6643
CP-LSH 0.7801 0.7128
N-LSH 0.4142 0.4307
@@
[…]
Ours-$3$ 0.9175 0.7879
Results and Analysis. We present performance curves across datasets, reporting Recall10@10 and Recall@1 as a function of the number of retrieved candidates. As shown in Fig. 3, our method consistently outperforms state-of-the-art baselines across both metrics. Notably, it exhibits a significant advantage in the low-probe regime: for example, on SIFT1M, our method achieves Recall@1 of $\sim 69\%$ using $\sim 3400$ candidates, compared to 61% for Faiss-IVF and 55% for Faiss-IVFPQFS using similar candidate counts. This efficiency gap is most evident below the 80% recall threshold, where traditional methods struggle to match the accuracy without a sharp rise in probe cost. Fig. 3(a) highlights the efficiency trade-off between training time and retrieval quality, measured via Recall@1-per-probe. All our variants lie in the top-left region of the plot, offering options ranging from ultra-fast lightweight models (e.g., Ours 1%) to more accurate, higher-capacity variants (e.g., Ours 10%). In contrast, competing methods tend to occupy lower-efficiency or higher-cost zones, demonstrating a less favorable balance between indexing overhead and retrieval quality.
(a) Generalization
(b) Time-recall@1
(c) Quantization
(d) $|\mathcal{G}|$ impact
Figure 4: (a) Generalization with limited training data. (b) Time-recall@1 tradeoff. (c) Quantization: efficient search via partial bucket scanning. (d) Impact of initial number of Gaussians.
Furthermore, for classification tasks on MNIST and Fashion-MNIST, our approach demonstrates superior classification accuracy compared to the other benchmarks, shown in Tab.1. This enhancement is consistent with theoretical expectations, as our methodology learns the inherent probabilistic structure of the training data rather than depending on predetermined heuristics, static partitioning, or hash-based approximations. By dynamically adapting to the statistical properties of the data distribution, our method identifies more semantically coherent neighbors. Our results suggest that GARLIC acts as a local manifold modeler, not just an indexing method offering smoother, distribution-aware retrieval and classification.
Table 2: Effect of loss terms and adaptive refinement on average Recall10@10 / Probe (×$\times 10^{5}$). Higher is better.
@@
Loss terms
   w/ all 10.43
   $\mathcal{L}_{div}+\mathcal{L}_{cov}$ 9.82
   $\mathcal{L}_{div}+\mathcal{L}_{anchor}$ 7.87
   $\mathcal{L}_{div}$ 7.17
[…]
   w/ both 10.43
   w/ split 8.65
   w/ clone 8.47
   w/o any 2.79
Ablation study. We investigate the robustness and data efficiency of our method through controlled downsampling of the training set. As shown in Fig. 4(a), our method consistently maintains strong performance across varying training sizes. Notably, even with only $1\%$ of the training data, our model outperforms the Faiss-IVF variants of up to $10\%$ training size in terms of Recall@1 versus the number of retrieved candidates. This trend remains consistent across higher percentages, demonstrating that our approach learns a compact yet highly effective representation of the data distribution. Furthermore, Fig. 4(b) highlights the favorable scaling of our method with respect to training time, while Fig. 4(c) demonstrates the impact of our quantization scheme. In particular, it exhibits strong performance across a broad quantization range ($20-100\%$), with only the most aggressive setting (10%) leading to degradation. This indicates that our quantization strategy is robust and well-aligned with our model structure. Fig. 4(c) shows that increasing the number of initial Gaussians accelerates convergence. However, the primary role of adaptive refinement is not to fix the final count, but to control their spatial density.
Tab. 2 analyzes the impact of loss terms and adaptive refinement. Split and clone are pivotal for performance, while the covariance term ($\mathcal{L}_{\text{cov}}$) significantly boosts retrieval. The anchor term ($\mathcal{L}_{\text{anchor}}$) has minor effect on raw performance but is essential for making Gaussians geometrically informative. Additional ablations and results, covering hyperparameter sensitivity, design variants, and the ability of Gaussians to capture non-linear, informative structure, are provided in the SM.
Limitations. GARLIC uses Mahalanobis distance, which assumes Euclidean structure and is not directly compatible with angular similarity. This can be addressed by replacing Gaussians with angular analogues such as von Mises–Fisher distributions or by embedding into hyperspherical spaces. Although GARLIC performs strongly in low-probe regimes, achieving very high recall (e.g., $\geq 95\%$) requires more computation. This recall-latency trade-off can be mitigated by increasing the number of Gaussians and structuring them with trees (e.g., KD-tree or Ball-tree over Gaussian means), reducing query cost. The space complexity scales quadratically with the dimension ($O(Kd^{2})$) due to full anisotropic covariances; however, this can be alleviated by using diagonal or low-rank approximations, or by quantizing the Cholesky factors. Lastly, while the method is mildly sensitive to initialization, its impact is reduced by the information-theoretic objectives and the progressive refinement through split and clone operations.
4 Conclusions
We introduced GARLIC, a geometric structure that learns the underlying distribution for both approximate nearest neighbor search and classification. By combining information-theoretic objectives with adaptive refinement (split and clone), and representing the space via anisotropic Gaussians with local quantization, our method achieves strong performance in both retrieval and classification, particularly in low-probe regimes. Experiments demonstrate competitive recall-efficiency tradeoffs and robustness under severe data reduction, highlighting its generalization capabilities.
Acknowledgments
PR and IE were partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program. PR, CT, IC, GI and IE have been partially supported by the European Union and the “Greece 2.0” national recovery and resilience plan, under the call "RESEARCH-CREATE-INNOVATE" (Code: ΤΑΕΔΚ-06168).
Appendix A Supplementary Material
This supplementary material provides additional details and experiments as mentioned in the main paper. Sec. A.1 discusses and provides detailed computational time and space complexity for the build and query procedures of our proposed method. Sec. A.2 contains additional technical details regarding the datasets used and the experimental setup, as well as the training pipeline. Sec. A.3 provides ablation studies to validate our choices regarding the tuned hyperparameters of our method and the components of it (discussed in Sec.4 of the main paper). Sec. A.4 contains visualization of diagnostics and statistics regarding our proposed method, in order to provide a comprehensive understanding of its behavior.
A.1 Complexity Analysis
Index build complexity.
Let $|X|$ be the number of data points, $d$ the embedding dimension, $K^{\prime}$ the initial number of Gaussians, and $r\ll d$ the reduced dimension used for local quantization. Let $I$ be the number of optimization steps. We denote by $S$, $C$, and $P$ the number of splits, clones, and pruned Gaussians, respectively, and define the final number of Gaussians as $K=K^{\prime}+S+C-P$. Let $|B_{g}|$ be the average bucket size, $c$ the number of K-Means clusters used during splitting, and $k^{\prime}$ the number of candidate points sampled per clone.
For the initialization, since we use K-Means++ on $K^{\prime}$ total cluster centers, we need $\mathcal{O}(K^{\prime}\cdot d\cdot|X|)$ time. For the optimization part, we need to perform a total number of $I$ iterations of full Mahalanobis-based point-to-Gaussian assignment, thus a total of $\mathcal{O}(I\cdot|X|\cdot K\cdot d^{2})$ worst-case time. Separate from the optimization, we analyze the split, clone and prune operations that are not applied on every iteration of the optimization. (i) The split operation runs DBSCAN or K-Means (with $c$ clusters), thus for a total of $S$ such operations we would need $\mathcal{O}(S\cdot(|B_{g}|\cdot d\cdot c+d^{2}))$; (ii) the clone operation locates the subset of points outside the Gaussian’s boundary (e.g., between $\tau\cdot\sigma$ and $(1+\epsilon)\cdot\tau\cdot\sigma$) for which it identifies new local modes, thus for a total of $C$ operations, it leads to $\mathcal{O}(C\cdot(k^{\prime 2}\cdot d+d^{2}))$; and (iii) the prune operation simply removes low-cardinality Gaussians and reassigns the points to the nearest active Gaussian, which takes $\mathcal{O}(P\cdot|B_{g}|\cdot d)$ time. For the quantization of each Gaussian, PCA is performed on all points inside the Gaussian, which projects the data into reduced local bases. In total, for the quantization we need $\mathcal{O}(K^{\prime}\cdot|B_{g}|\cdot d^{2}+|X|\cdot d\cdot r)$ time.
Query complexity.
Space complexity.
Let $K$ be the number of Gaussians, $d$ the data dimension, and $N=|X|$ the dataset size. The model stores mean vectors $\boldsymbol{\mu}\in\mathbb{R}^{K\times d}$, Cholesky parameters $\mathbf{L}\in\mathbb{R}^{K\times d\times d}$, and Buckets storing point indices, requiring $\mathcal{O}(N)$ space. Thus, the total space complexity is:
$\mathcal{O}\left(K\cdot(d^{2}+d)+N\right)$
@@ A.2 Implementation details
We evaluate our method on three standard benchmark datasets: (1) SIFT1M [37], containing one million 128-dimensional SIFT descriptors that capture scale- and rotation- invariant local image features; (2) MNIST [32], consisting of 70,000 grayscale handwritten digit images (28×28 pixels) flattened to 784-dimensional vectors; and (3) Fashion-MNIST [44], a more challenging variant with the same format but featuring 10 categories of fashion items. For retrieval tasks, we used the ANN-Benchmark [8] versions of these datasets (available at http://ann-benchmarks.com/sift-128-euclidean.hdf5, http://ann-benchmarks.com/mnist-784-euclidean.hdf5, and http://ann-benchmarks.com/fashion-mnist-784-euclidean.hdf5) to ensure a standardized comparison with existing methods. For MNIST and Fashion-MNIST classification experiments, we utilized the PyTorch torchvision implementations to maintain consistency with established classification benchmarks.
@@
Our method was implemented in Python, using optimized libraries such as PyTorch [40] and NumPy [22]. The experiments were carried out on an Intel Core i7-7820X CPU (16 threads), a Quadro RTX 8000 GPU (48 GB VRAM) and 125 GiB of RAM.
@@
Batch sizes of $20000$ are used for SIFT, and $5000$ for MNIST and Fashion-MNIST datasets, with normalization by subtracting the mean and dividing by the standard deviation. The number of training epochs is $250$, with a typical early stop at $120$. Gaussian updates are scheduled with two phases: a warm-up phase lasting $35$ epochs and an optimization phase where structural operations like splitting and cloning are triggered every $35$ epochs and pruning every $60$.
@@
Splitting is applied to Gaussians with cardinality exceeding a fraction of the dataset $\gamma=1\times 10^{-2}$, using DBSCAN or K-Means, as fallback, with $c=2$ clusters. The covariance of each new Gaussian is scaled down by $9\times 10^{-1}$. Cloning selects dense regions just outside the Gaussian boundary, defined by a Mahalanobis shell with inner threshold $\tau$ and outer threshold $(1+\epsilon)\cdot\tau$, where $\epsilon=2.2$ controls the shell thickness. From the set of points referenced in $(\tau,e\tau]$, a random selection of $60\%$ is made. Cloning is not performed on a Gaussian unless its cardinality surpasses a threshold specified by $8\times 10^{-4}|X|$. Gaussians that have degenerated into a single point are eliminated. Pruning is executed at intervals of every $60$ epochs.
@@
Local PCA is performed per bucket using top-$3$ eigenvectors. Reduced points are quantized using a spherical grid with $n_{\text{radial}}=6$ and $n_{\text{angular}}=4$, forming directional bins per Gaussian.

### v2

@@ 2.5 Inference
3 Experiments
Datasets. We evaluate our method on three benchmark datasets: SIFT1M (Lowe, 2004) (128-dimensional image descriptors with one million points), MNIST (LeCun et al., 1998) and Fashion-MNIST (Xiao et al., 2017) (784-dimensional vectors from $28\times 28$ grayscale images) following the standard setup in ANN-Benchmarks (Aumüller et al., 2020). Further information about the datasets can be found in the Appendix.
Evaluation. We evaluate GARLIC on the approximate nearest neighbor task, reporting Recall@1 and as function of the number of distance computations while varying the probe budget. Recall@1 measures the accuracy of the very first retrieved neighbor, reflecting ranking performance, and is strict. In Appendix A.1, we showcase more related metrics such as $\epsilon$-Recall, empricial c-approximation factors and mean relative error. We use distance computations, rather than wall-clock latency (QPS), as the efficiency axis. Distance computations capture the algorithmic effort of a search procedure independently of hardware and implementation details (different CPUs/GPUs, BLAS kernels, batching, I/O) (Aumüller et al., 2020, Peng et al., 2023). This metric directly reflects the goal of space-partitioning methods, which is to minimize the number of distances required to achieve a target recall.
(a) Generalization
(b) Batch size
(c) $|\mathcal{G}|$ impact
(d) Initialization
(e) Split threshold $\gamma$
(f) Clone threshold $e$
[…]
Baselines and Comparisons. We evaluate GARLIC against representative approximate nearest neighbor methods such as $k$-Means++ (Arthur & Vassilvitskii, 2006), BIRCH (Zhang et al., 1996), Gaussian Mixture Models (GMM) (Dempster et al., 1977, Banerjee et al., 2005), Neural-LSH (N-LSH) (Dong et al., 2020), Hyperplane LSH (HP-LSH) (Indyk & Motwani, 1998), Cross-polytope LSH (CP-LSH) (Andoni et al., 2015), ITQ (Gong et al., 2012), PCA Tree (Kumar et al., 2008, Abdullah et al., 2014, Sproull, 1991, McNames, 2001), Faiss-IVF and Faiss-IVFPQFS (Douze et al., 2024). All baselines were run under identical conditions, using the standardized ANN-Benchmarks datasets and query splits, with hyperparameters optimized according to their respective publications and usage.
Results and Analysis. We present performance curves across datasets, reporting Recall@1 as a function of the number of retrieved candidates. As shown in Figure 3, our method performrs efficiently with respect to other baselines. GARLIC generally achieves preferable recall–efficiency trade-offs than traditional space partitioning methods, including inverted-file indices (Faiss-IVF, IVFPQFS), hashing approaches (HP-LSH, CP-LSH, Neural-LSH), and tree-based methods (PCA-Tree, BIRCH). Figure 3(a) evaluates efficiency through Recall@1-per-probe versus build time. All GARLIC variants occupy the upper-left region, indicating favorable trade-offs: lightweight models trained with $1\%$ of the data already surpass strong baselines, while larger-capacity variants (e.g., $10\%$ training) further improve recall without excessive cost. In contrast, competing methods cluster in lower-efficiency or higher-cost regions, reflecting less balanced trade-offs between indexing overhead and search quality.
Ablation study. We conduct ablation studies to isolate the effect of key design choices and components. Unless stated otherwise, experiments are on SIFT-1M with a training subset of $5\%$ of $|X|$, and Recall@1 is measured against distance computations. We investigate the robustness and data efficiency of our method through controlled downsampling of the training set. As shown in Figure 4(4(a)), our method consistently maintains strong performance across varying training sizes. Notably, even with only $1\%$ of the training data, our model outperforms the Faiss-IVF variants of up to $10\%$ training size in terms of Recall@1 versus the number of retrieved candidates. This trend remains consistent across higher percentages, demonstrating that our approach learns a compact yet highly effective representation of the data distribution. Furthermore, Figure 4(4(b)) highlights the batch-size effect (here using a fixed training subset of $100k$ points instead of $50k$), where larger batches yield consistently better performance. Finally, Figure 4(4(c)) shows that more initial Gaussians produce better results at equal iterations, reflecting faster convergence. This happens because adaptive density control targets spatial placement rather than Gaussian count.
Table 1: Effect of loss terms and adaptive refinement average Recall@1 / Candidates ($\times 10^{5}$). Higher is better.
Configuration Performance $\uparrow$
Loss terms
   w/ all 16.20
   $\mathcal{L}_{div}+\mathcal{L}_{cov}$ 15.38
   $\mathcal{L}_{div}+\mathcal{L}_{anchor}$ 10.64
   $\mathcal{L}_{div}$ 19.78
[…]
   w/ both 16.20
   w/ split 14.22
   w/ clone 15.45
   w/o any 14.63
We next evaluate core design hyperparameters. In Figure 4(4(d)), we investigate initialization strategies for Gaussian centers, revealing that K-Means initialization consistently outperforms alternatives. Although farthest point sampling (FPS) achieves comparable recall, it requires approximately 50% more probes, and catching up to K-Means requires more training time. In Figure 4(4(e)), the split threshold $\gamma$, controlling the density condition for Gaussian subdivision, is examined. The optimal values are $\gamma=0.005$ and $\gamma=0.01$, though $\gamma=0.005$ requires more time due to more frequent splits. As for the clone threshold $e$, depicted in Figure 4(4(f)), which determines the outer shell for cloning, it remains stable across different values, with $e=2.2$ being preferable in situations requiring high recall.
Table 1 analyzes the impact of loss terms, adaptive refinement and covariance configurations. Split and clone are pivotal for performance, while the covariance term ($\mathcal{L}_{\text{cov}}$) significantly boosts retrieval. The anchor term ($\mathcal{L}_{\text{anchor}}$) has minor effect on raw performance but is essential for making Gaussians geometrically informative. Additional ablations and results are provided in the Appendix.
Limitations and Future Work. GARLIC employs Mahalanobis distance and Gaussian primitives, which assume a Euclidean metric and are therefore not directly compatible with angular similarity. This may be addressed by adopting angular counterparts, such as von Mises–Fisher distributions, which we leave for future work. The trade-off between recall and latency can be further enhanced by augmenting the number of Gaussians and organizing them within tree structures (e.g., KD-tree or Ball-tree over Gaussian means), reducing query cost. The spatial complexity scales quadratically with dimensionality ($\mathcal{O}(Kd^{2})$) due to full anisotropic covariances; this challenge can be mitigated through low-rank approximations or quantization of Cholesky factors. Finally, while the method shows some sensitivity to initialization, this is alleviated by the information-theoretic objectives and the progressive refinement via split and cloning. Beyond these limitations, GARLIC is naturally incremental: each Gaussian cell maintains sufficient statistics (mean, covariance, cardinality), allowing new data to be integrated through online updates. Combined with local split/clone refinement, this enables streaming and online learning scenarios without rebuilding the entire index. We leave a systematic evaluation of this capability to future work.
4 Conclusions
We introduced GARLIC, a geometric structure that learns the underlying distribution for both approximate nearest neighbor search and classification. By combining information-theoretic objectives with adaptive refinement (split and clone), and representing the space via anisotropic Gaussians, our method achieves competitive performance in Euclidean approximate nearest neighbor, particularly in low-probe regimes. Experiments demonstrate competitive recall-efficiency tradeoffs and robustness under severe data reduction, highlighting its generalization capabilities.
Appendix A Appendix
This appendix provides additional details and experiments as mentioned in the main paper. Section A.1 provides additional results and ablation studies to the ones presented in Section 3 of the main paper. Section A.1.1 contains visualization of diagnostics and statistics regarding our proposed method, to provide a comprehensive understanding of its behavior. Section A.2 contains additional technical details regarding the datasets used and the experimental setup, as well as the training pipeline. Section A.3 discusses and provides detailed computational time and space complexity for the build and query procedures of GARLIC.
A.1 Further Results
[…]
OursK-MeansBIRCHCP-LSH
HP-LSHPCA TreeFaiss-IVFFaiss-IVFPQFS
[…]
Refer to captionMNIST
Refer to captionFMNIST
[…]
Refer to captionPrecision @ 10
[…]
We evaluate retrieval quality using both accuracy- and distortion-based criteria, as shown in Figures 5 and 6. Beyond Recall@$1$, we report Precision@10 (P@10), capturing the fraction of retrieved points among the top-10 that are true neighbors. To assess approximation tightness for the nearest neighbor, we measure the distance ratio between the returned neighbor and the exact nearest neighbor. From this we derive: (i) $\varepsilon$-Recall, the fraction of queries where the retrieved distance is within a factor $(1+\varepsilon)$ of the ground-truth (with $\varepsilon\in\{0.01,0.10\}$); (ii) the percentiles $r95$ and $r99$, which correspond to empirical $c$-approximation factors at the 95th and 99th percentiles; and (iii) the mean relative error, $RE_{\text{mean}}$, summarizing the average stretch beyond the true nearest-neighbor distance. These metrics complement recall by quantifying how close the returned distances are to the exact nearest neighbor, and follow standard practice in ANN-Benchmarks.
OursK-MeansBIRCHCP-LSH
HP-LSHPCA TreeFaiss-IVFFaiss-IVFPQFS
Refer to captionr95SIFT
Refer to captionMNIST
Refer to captionFMNIST
Refer to captionr99
@@ A.1 Further Results
Across datasets we observe consistent trends: $\varepsilon$-Recall at $1\%$ and $10\%$ rapidly approaches $1.0$, indicating that even under small probe budgets the vast majority of retrieved neighbors lie within $1\%$$10\%$ of the true nearest-neighbor distance. The $r95$ and $r99$ curves remain close to $1.0$, showing that $5\%99$$9\%$ of queries admit near-exact retrieval, while $RE_{\text{mean}}$ stays low and stable, confirming that average distortion is minimal. Together with high P@10, these results demonstrate that the learned Gaussian partitions not only achieve strong recall but also return candidates that are quantitatively close to the exact nearest neighbors, providing both efficiency and fidelity in the ANN process.
Additional Abblation Studies
In Figure 7(7(a)), the contribution of each loss term to the optimization process is analyzed. The results indicate that the covariance loss $\mathcal{L}_{\text{cov}}$ produces the most significant performance improvement, while the anchor loss $\mathcal{L}_{\text{anchor}}$ functions as an effective regularizer, grounding each Gaussian by aligning it with its corresponding point distribution. Figure 7(7(b)) demonstrates the effects of our adaptive refinement operations, where the split and clone mechanisms improve retrieval performance when used together, each contributing by providing complementary benefits to the quality of representation. Figure 7(7(c)), addresses the parameter $\tau$ within $\mathcal{L}_{div}$, demonstrating that $\tau=3$ is the optimal choice, particularly in scenarios involving a limited number of probes. Values surpassing those presented in our study ($\tau>3$) result in less optimal outcomes. For instance, $\tau=4$ demonstrates a Recall@1 value of 0.67~0.67 when evaluated with $154264$ probes.
(a) Loss terms
(b) Split & clone impact
(c) ReLU parameters $\tau$
(d) Quantization
(e) Dimensionality effect
@@
Furthermore, Figure 7(7(d)) demonstrates the impact of our quantization scheme. In particular, it exhibits strong performance across a broad quantization range ($20{-}100\%$), with only the most aggressive setting (10%) leading to degradation. This indicates that our quantization strategy is robust and well-aligned with our model structure. Finally, Figure 7(7(e)) examines dimensionality effects using Fashion-MNIST data resized to various dimensions (to simulate higher-dimensional embeddings). The results confirm that GARLIC maintains strong performance across a wide range of dimensionalities.
Table 2: Effect of anisotropy on average Recall@1 / Probe ($\times 10^{5}$). Higher is better. As anisotropy decreases, performance degrades due to excessive probe usage.
@@
Covariance structure
   Anisotropic 16.20
   Diagonal 11.22
   Isotropic 10.75
Covariance structure vs performance.
We conduct a targeted study to assess the impact of different covariance configurations-namely, full (anisotropic), diagonal, and isotropic-affect the performance of GARLIC. As shown in Table 2, reducing the Gaussian expressiveness from full to diagonal and then to isotropic leads to a notable decline in average Recall@1 per probe. While diagonal and isotropic configurations reduce both the parameter count and computation per Gaussian to a linear level, $\mathcal{O}(k\cdot d)$, they lead to aggressive pruning and an increased number of probes to make up for the representational loss. This suggests that space complexity cannot be drastically reduced without harming retrieval quality because simpler Gaussian parameterizations result in degraded locality and coverage.
A.1.1 Qualitative Results
Coverage and reconstruction diagnostics.
Curvature-based diagnostics.
Figure 9 provides four views exploring the relationship between curvature and structural properties of the Gaussian assignments across datasets, to examine wheather the learned model is informative and geometrically consistent with the data $x\in X$.
Figure 10 demonstrates the capability of GARLIC to capture the intrinsic geometric structure of high-dimensional data with locally varying dimensionality. The anisotropy histograms (first row) reveal how Gaussians adapt to regions of different local dimensionality, with $\lambda_{\text{max}}/\lambda_{\text{min}}$ ratios ranging from nearly isotropic to highly stretched configurations across all datasets. Rather than simply partitioning space uniformly, Gaussians adapt their shapes to the underlying manifold structure, as confirmed by the diagonality histograms (third row) and effective dimension measurements (fourth row). This representation allows us to model stratified data where different intrinsic dimensionalities coexist, allowing the data to become pancake-like for surface regions, needle-like for curve regions, and ball-like for volumetric areas. Unlike traditional space partitioning methods, GARLIC models the underlying data distribution probabilistically, not just approximating distances for retrieval. The explained variance surfaces (bottom row) show that Gaussians efficiently capture the relevant dimensions at each location, enabling estimation of true manifold distances rather than just Euclidean distances to samples. This provides more semantically meaningful results in regions where the intrinsic dimensionality is lower than the ambient space. This completes the supplementary material.
@@ A.2 Implementation details
We evaluate our method on three standard benchmark datasets: (1) SIFT1M Lowe (2004), containing one million 128-dimensional SIFT descriptors that capture scale- and rotation- invariant local image features; (2) MNIST LeCun et al. (1998), consisting of 70,000 grayscale handwritten digit images (28×28 pixels) flattened to 784-dimensional vectors; and (3) Fashion-MNIST Xiao et al. (2017), a more challenging variant with the same format but featuring 10 categories of fashion items. For retrieval tasks, we used the ANN-Benchmark Aumüller et al. (2020) versions of these datasets (available at http://ann-benchmarks.com/sift-128-euclidean.hdf5, http://ann-benchmarks.com/mnist-784-euclidean.hdf5, and http://ann-benchmarks.com/fashion-mnist-784-euclidean.hdf5) to ensure a standardized comparison with existing methods.
@@
Our method was implemented in Python, using optimized libraries such as PyTorch Paszke (2019) and NumPy Harris et al. (2020). The experiments were carried out on an Intel Core i7-7820X CPU (16 threads), a Quadro RTX 8000 GPU (48 GB VRAM) and 125 GiB of RAM.
@@
Batch sizes of $50000$ are used for SIFT, and $20000$ for MNIST and Fashion-MNIST datasets, with z-score normalization by subtracting the mean and dividing by the standard deviation. The number of training epochs is $250$, with a typical early stop at $120$. Gaussian updates are scheduled with two phases: a warm-up phase lasting $35$ epochs and an optimization phase where structural operations like splitting and cloning are triggered every $35$ epochs and pruning every $60$.
@@
Splitting is applied to Gaussians with cardinality exceeding a fraction of the dataset $\gamma=1\times 10^{-2}$, using DBSCAN or K-Means, as fallback, with $c=2$ clusters. The covariance of each new Gaussian is scaled down by $9\times 10^{-1}$. Cloning selects dense regions just outside the Gaussian boundary, defined by a Mahalanobis shell with inner threshold $\tau$ and outer threshold $(1+e)\cdot\tau$, where $e=2.2$ controls the shell thickness. From the set of points referenced in $(\tau,e\tau]$, a random selection of $60\%$ is made. Cloning is not performed on a Gaussian unless its cardinality surpasses a threshold specified by $8\times 10^{-4}|X|$. Gaussians that have degenerated into a single point are eliminated. Pruning is executed at intervals of every $60$ epochs.
@@
Local PCA is performed per cell using top-$3$ eigenvectors. Reduced points are quantized using a spherical grid with $n_{\text{radial}}=6$ and $n_{\text{angular}}=4$, forming directional bins per Gaussian.